OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3).

OpenStudy (anonymous):

Perpendicular lines have negative reciprocal slopes.

OpenStudy (anonymous):

do you understand what Ruka said ^ ?

OpenStudy (anonymous):

Once you find the slope, sub the values of x and y into the equation to find the "equation of a line" to get your answer

OpenStudy (anonymous):

>.> uh..

OpenStudy (anonymous):

For example: y=mx+b the slope is m. For the perpendicular line, the slope will be -1/m

OpenStudy (anonymous):

do you know what perpendicular means?

OpenStudy (anonymous):

Plug the values in for y-y1=m(x-X1) once you have your slope.

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

perpendicular means a line forms a right angle with another line... the angle between them is 90 degrees.

OpenStudy (anonymous):

Oh.

OpenStudy (anonymous):

so in terms of slope, two lines that are perpendicular to each other have negative reciprocal slopes... for example, what's the slope of y = 2x + 2?

OpenStudy (anonymous):

-2?

OpenStudy (anonymous):

no. m=2, the slope of y = 2x + 2 is 2

OpenStudy (anonymous):

so the negative reciprocal of that is -1/m or -1/2. do you understand?

OpenStudy (anonymous):

Kind of.

OpenStudy (anonymous):

So much to remember in math.

OpenStudy (anonymous):